THE X^{2}Π_{i}, A^{2}Δ_{i} AND B^{2}Σ^{+} LOW-LYING STATES OF NiCl : LASER INDUCED AND FOURIER TRANS-FORM EXPERIMENTS

Y. KROUTI, Laboratoire de Physique atomique et moléculaire, Faculté des Sciences Ain Chock B. P. 5366 Maârif, Casablanca, MOROCCO; A. POCLET, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523 Centre d’Etudes et de Recherches Laser et Applications, Université de Lille I59655, Villeneuve d’Ascq Cedex, FRANCE; \underline{T. HIRAO}, Department of Chemistry, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1; B. PINCHEMEL, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523 Centre d’Etudes et de Recherches Laser et Applications, Université de Lille I59655, Villeneuve d’Ascq Cedex, FRANCE; and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, CANADA N2L 3G1.

The X^{2}Π_{3/2}, A^{2}Δ_{5/2} (161 cm^{−1}) and B^{2}Σ^{+} (1,768 cm^{−1}) low-lying states of NiCl have been recently identified. Transitions linking these states to upper [21.9]^{2}Δ_{5/2}, [24.5]^{2}Π_{3/2} and [24.9]^{2}Π_{1/2} states have been rotationally analyzed. Laser induced fluorescence experiments allowed the identification of the X^{2}Π_{1/2} (382 cm^{−1}) and A^{2}Δ_{3/2} (1,646 cm^{−1}) spin-orbit components. All the low-lying states of NiCl associated with the nickel 3d atomic orbital are now observed. In addition we identified a new electronic state located at 21,608 cm^{−1}. This state is linked to all the five spin-orbit components of the X , A and B states. High resolution Fourier transform experiments made possible the analysis of three of these transitions. Laser induced dispersed fluorescence suggest that this state is also linked to a new low-lying state (1,378 cm^{−1}) which could be a component of a quartet state of NiCl.